OpenStudy (anonymous):

Solve the following system of equations and show all of your work. 4x + 3y = -1 3x + y = 3

OpenStudy (tyteen4a03):

4x + 3y = 1 3x + y = 3 Notice the bare y in the second equation. In this case, make use of this to derive a third equation from the second: y = 3 - 3x. Now substitute this into the first equation and find x. Finally use the newfound value of x to find y.